OpenStudy (anonymous):

Which fraction is equivalent to 9/16?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What fractions do you have to choose from?

OpenStudy (anonymous):

A. 12/48 B. 18/48 C. 45/80 D. 56/96

jimthompson5910 (jim_thompson5910):

ok, let's check to see if 9/16 = 12/48 is true

jimthompson5910 (jim_thompson5910):

Start with 9/16 = 12/48 and cross multiply to get 9/16 = 12/48 9*48 = 16*12 what do you get when you multiply out the two sides?

OpenStudy (anonymous):

1.333333333

OpenStudy (anonymous):

I feel like I did it wrong

jimthompson5910 (jim_thompson5910):

not sure how you got that

OpenStudy (anonymous):

ok im gonna do step by step lol

OpenStudy (anonymous):

9/16 x 12/48 is 9/64

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):